# Family Unification in $SO(16)$ Grand Unification

### v1

@@ 1 Introduction
The existence of three chiral generations of quarks and leptons is one of the most mysterious fact in particle physics. In addition, the mass hierarchy among different flavors of fermions generated by the Higgs mechanism via their corresponding Yukawa couplings strongly calls for a better understanding of the underlying structure of Nature. There are many attempts to understand the origin of chiral generations and/or their mass hierarchy by considering, e.g., the so-called horizontal symmetry (or family symmetry) in four-dimensional (4D) theories [1, 2, 3, 4, 5, 6, 7, 8, 9], as well as employing special geometrical structures in higher-dimensional theories [10, 11].
@@
There are some attempts to unify GUT and family groups into a larger GUT group in 4D and higher-dimensional theories [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. In particular, models based on $SU(19)$ GUT gauge symmetry are discussed in the 4D framework [32, 37] and the six-dimensional (6D) framework by one of the authors [36]. In these models, the three chiral generations of the SM fermions are unified into a fermion multiplet in the second-rank antisymmetric tensor representation ${\bf 171}$ of $SU(19)$. In such models, $SU(19)$ is broken to a maximal regular subgroup $SU(16)\times SU(3)\times U(1)$, and $SU(16)$ is further broken to a maximal special subgroup $SO(10)$. The branching rule of ${\bf 171}$ under $SU(19)\supset SO(10)\times SU(3)\times U(1)$ is given by [38]
@@
(For more information about Lie algebras and their subalgebras, see, e.g, Ref. [38].)
In this paper, we will propose a new model based on $SO(16)$ GUT gauge symmetry that unifies the three chiral generations of SM fermions into a 6D Weyl fermion in a spinor representation of $SO(16)$ without exotic fermions on the 6D spacetime $M^{4}\times D^{2}/\mathbb{Z}_{N}$ $(N\geq 10)$. The 6D $SO(16)$ gauge anomaly is canceled by the vectorlike nature of the model, and the 4D gauge anomalies are canceled by introducing suitable 4D localized fermions at the fixed point. The gauge coupling constant of $SO(16)$ is asymptotically free.
@@
Let’s consider the 6D spacetime $M^{4}\times D^{2}/\mathbb{Z}_{N}$ with Cartesian coordinates $x^{\mu}$ ($\mu=0,1,2,3$) and circular coordinates $(r,\theta)$ or a complex coordinate $z=re^{i\theta}$, where $D^{2}$ is a 2-dimensional (2D) real manifold and there is a fixed point at $r=0$ [41]. We take the outer radius of the disc $D^{2}$ is $R$. A $\mathbb{Z}_{N}$ symmetry on $D^{2}$ is defined as
@@
In the following, we discuss the gauge fields associated with the $SO(16)$ generators $T^{\alpha}$ ($\alpha=1,2,...,120$), which are expressed using the Pauli matrices $\sigma_{a}$ ($a=1,2,3$) and $16\times 16$ symmetric and anti-symmetric matrices $S_{8}$ and $A_{8}$ as [42]:
@@
where $\bar{\omega}:=\omega^{-1}$, $\eta_{\Phi}\subset\mathbb{Z}_{N}$, and $\ell_{\Omega}$ and $m_{\Omega}$ are non-negative integers depending on the representation of $G$. (In detail, see, e.g., Refs. [28, 41].) We will apply this discussion for the $SO(16)$ gauge field $A_{\mu}$ in the 4D spacetime below, noting that different choices of $R_{\Omega}$ will lead to different symmetry breaking schemes.
@@
This symmetry breaking pattern is realized by taking the following matrix:
$R_{\Omega}=\left(\begin{array}[]{cc}0&\omega^{-a}I_{8}\\ \omega^{+a}I_{8}&0\\ \end{array}\right),$ (2.12)
where $a=1,2,\cdots,N-1$ (mod. $N$), $N$ must be even to satisfy $R_{\Omega}^{N}=I_{16}$, and the exponents of $\omega$ are proportional to the $U(1)_{a}$ charges. In this case, the gauge field $A_{\mu}$ is transformed under $\mathbb{Z}_{N}$ as
@@
$R_{\Omega}=\left(\begin{array}[]{cccc}0&0&\omega^{-a-3b}I_{5}&0\\ 0&0&0&\omega^{-a+5b}I_{3}\\ \omega^{+a+3b}I_{5}&0&0&0\\ 0&\omega^{+a-5b}I_{3}&0&0\\ \end{array}\right),$ (2.23)
where $|a+3b|,|a-5b|\in\mathbb{Z}$, and the exponents of $\omega$ are proportional to the $U(1)_{a}$ and $U(1)_{b}$ charges.
@@
${\bf 120}$ $=({\bf 45,1})(0)\oplus({\bf 1,8})(0)\oplus({\bf 1,1})(0)\oplus({\bf 1,3})(4)\oplus({\bf 1,\overline{3}})(-4).\oplus({\bf 10,3})(4)\oplus({\bf 10,\overline{3}})(-4).$ (2.24)
This symmetry breaking pattern is realized by choosing $a=-3\ell/4$ and $b=\ell/4$ $(\ell\in\mathbb{Z})$ in Eq. (2.23) as
$R_{\Omega}=\left(\begin{array}[]{cccc}0&0&I_{5}&0\\ 0&0&0&\omega^{+2\ell}I_{3}\\ I_{5}&0&0&0\\ 0&\omega^{-2\ell}I_{3}&0&0\\ \end{array}\right),$ (2.29)
where the exponents of $\omega$ are proportional to the $U(1)$ charges. This matrix breaks $SO(16)$ to $SO(10)\times SU(3)\times U(1)$. Note that if we choose special values of $N$ and $\ell$, then $SO(16)$ is unbroken for $\ell=0$ (mod. $N/2$), and $SO(16)$ is broken to $SO(10)\times SO(6)\simeq SO(10)\times SU(4)$ for $\ell=N/4$ (mod. $N/2$), where $\ell\in\mathbb{Z}$. We will consider the $SO(16)$ breaking to $SO(10)\times SU(3)\times U(1)$ in the following sections.
@@ 3 Gauge symmetry and gauge boson
From Table 1, we find that there are the following symmetry breaking patterns, depending on the value of $\ell$ for $N\not=|2\ell|,|4\ell|$:
@@
We focus on the symmetry breaking pattern $SO(16)\to SO(10)\times SU(3)\times U(1)$ in Eq. (3.5), where the three chiral generations of SM fermions can be unified into one fermion multiplet. In the case, we can take $\ell\not=0$ (mod. $N/4$). Here, we further impose the condition that no extra zero mode appears from the extra dimensional components of the gauge field. From Table 1, we find the following constraints for $\ell$ and $N$ to avoid the unwanted zero modes:
@@
This leads to the requirement that $N\geq 6$ $(N\in\mathbb{Z}_{\rm even})$. For $\ell=2$, from Eq. (3.6), we find
@@
This leads to $N=6$ or $N\geq 10$ $(N\in\mathbb{Z}_{\rm even})$. We can also consider $\ell\geq 3$ and negative integers in the same manner.
@@ 4 Unification of the three chiral generations of SM fermions
$\Omega\Psi_{\bf 128}^{\alpha}(x^{\mu},z,\bar{z})=\Psi_{\bf 128}^{\alpha}(x^{\mu},\omega z,\bar{\omega}\bar{z})=\eta_{\Psi^{\alpha}}R_{\Omega}^{\rm sp}\Psi_{\bf 128}^{\alpha}(x^{\mu},\omega z,\bar{\omega}\bar{z}),$ (4.4)
@@
We need to find a set of $\ell$, $k$ and $N$ that simultaneously satisfy Eqs. (4.5) and (4.6). There turn out to be an infinite number of allowed sets of $\ell$, $k$ and $N$. Hence, we will only consider a few simple examples. First, we set $\ell=1$. In this case, Eq. (4.5) gives
@@
This leads to the requirement $N\geq 9$. Since $N$ must be even from $R_{\Omega}^{N}=I$, we conclude that $N\geq 10$ $(N\in\mathbb{Z}_{\rm even})$ in this case. In other words, we have now found a possible set of parameters, $(\ell,k,N)=(1,-4,10)$.
@@
This then leads to $N\geq 8$. Combining the two conditions, we need to take $N\geq 10$ $(N\in\mathbb{Z}_{\rm even})$. Thus, we have found another possible set of parameters: $(\ell,k,N)=(2,0,10)$.
In summary, to unify the three chiral generations of SM fermions without the occurrence of exotic fermions, one should take $N\geq 10$ $(N\in\mathbb{Z}_{\rm even})$ and $(\eta_{\Psi^{1}},\eta_{\Psi^{2}})=(\omega^{-1},\omega^{-4})$, $(\omega^{-2},1)$, $\dots$ In such cases, only the fermions in the $({\bf 16,3})(1)$ representation of $SO(10)\times SU(3)\times U(1)$ have zero modes.
@@
From the above discussions, it is clear that one choice of having the three chiral generations of SM fermions is to take $N\geq 10$ $(N\in\mathbb{Z}_{\rm even})$ and $(\eta_{\Psi^{1}},\eta_{\Psi^{2}})=(\omega^{-1},\omega^{-4})$. In this case, we obtain the 4D left-handed fermions in $({\bf 16,3})(1)$ of $SO(10)\times SU(3)\times U(1)$ from a single 6D fermion $\Psi_{\bf 128}^{\alpha=1}$ without having exotic fermions. In the following, we discuss 6D and 4D gauge anomaly cancellation in this scenario.
@@
$A_{\rm bulk}^{U(1)\mbox{-}U(1)\mbox{-}U(1)}=16\times 3\times(+1)^{3}=+16,$
@@
$A_{\rm bulk}^{\mbox{grav}\mbox{-}\mbox{grav}\mbox{-}U(1)}=16\times 3\times T_{SO(4)}{(\bf 2)}\times(+1)=+24,$ (5.1)
@@
We now examine what kind of 4D fields should be introduced to cancel out the gauge anomaly generated by the 6D fields given in Eq. (5.1). It turns out to be impossible to uniquely determine the required set of 4D fields solely by the requirement that the 4D gauge anomalies arising from the 6D fields cancel out. In the following, we provide an explicit example. One way is to introduce 4D localized left-handed fermions in $({\bf 16,3})(0)$, $({\bf\overline{16},\overline{3}})(-1)$, and $({\bf 1,\overline{3}})(0)$ of $SO(10)\times SU(3)\times U(1)$: $\psi_{L({\bf 16,3})(0)}$, $\psi_{L({\bf\overline{16},\overline{3}})(-1)}$, and $\psi_{L({\bf 1,\overline{3}})(0)}^{(\beta)}$ $(\beta=1,2,...,16)$. The gauge anomalies generated by the 6D fermions except for the $SU(3)\mbox{-}SU(3)\mbox{-}SU(3)$ anomaly in Eq. (5.1) are canceled by $\psi_{L({\bf 16,3})(0)}$ and $\psi_{L({\bf\overline{16},\overline{3}})(-1)}$. On the other hand, the $SU(3)\mbox{-}SU(3)\mbox{-}SU(3)$ anomaly is canceled by $\psi_{L({\bf 1,\overline{3}})(0)}^{(\beta)}$. When we take into account the vacuum expectation values (VEVs) of the $SU(3)$- and $U(1)$-breaking scalar fields, these 4D fields obtain masses that do not appear at low energies and will not be discussed further in this paper.
@@ 6 Symmetry breaking via the VEVs of scalars
Next, we consider the scalar content to realize the SM gauge symmetry at low energies. The set of scalar fields, which properly break the $SO(16)$ symmetry into $G_{\rm SM}$, is also not uniquely determined. Here we provide a working example by introducing localized scalars $\phi_{{\bf(10,1)}(0)}$, $\phi_{{\bf(45,1})(0)}$, $\phi_{{\bf(16,1)}(0)}$, $\phi_{{\bf(1,3)}(0)}$, $\phi_{{\bf(1,6)}(0)}$, $\phi_{{\bf(1,1)}(+1)}$ to realize $G_{\rm SM}$ at low energies. In this scenario, the symmetry breaking pattern is given as follows:
$\begin{array}[]{rll}SO(16)\ \to&SO(10)\times SU(3)\times U(1)&\mbox{by the non-trivial space structure}\\ \to&SO(10)&\mbox{by the VEVs}\ \langle\phi_{{\bf(1,3)}(0)}\rangle,\langle\phi_{{\bf(1,6)}(0)}\rangle,\langle\phi_{{\bf(1,1)}(1)}\rangle\not=0\\ \to&G_{\rm SM}&\mbox{by the VEVs}\ \langle\phi_{{\bf(45,1)}(0)}\rangle,\langle\phi_{{\bf(16,1)}(0)}\rangle\not=0\\ \to&SU(3)_{C}\times U(1)_{\rm EM}&\mbox{by the VEV}\ \langle\phi_{{\bf(10,1)}(0)}\rangle\not=0,\\ \end{array}$ (6.5)
where we assume that the Kaluza-Klein (KK) mass scales $1/R$ and the nonvanishing VEVs are at the order of $M_{\rm GUT}\simeq O(10^{16}-10^{17})$ GeV. By assuming symmetry breaking with the scalar field VEVs given above, all the introduced 4D fermions will attain large mass and will not appear at low energies. The $SO(10)$ GUT symmetry breaking to the SM gauge symmetry has been thoroughly examined in, for example, Refs. [46, 47, 48, 49, 38] and will not be discussed further here.

### v2

# YITP-25-45

@@ 1 Introduction
The existence of three chiral generations of quarks and leptons is one of the most mysterious facts in particle physics. In addition, the mass hierarchy among different flavors of fermions generated by the Higgs mechanism via their corresponding Yukawa couplings strongly calls for a better understanding of the underlying structure of Nature. There are many attempts to understand the origin of chiral generations and/or their mass hierarchy by considering, e.g., the so-called horizontal symmetry (or family symmetry) in four-dimensional (4D) theories [1, 2, 3, 4, 5, 6, 7, 8, 9], as well as employing special geometrical structures in higher-dimensional theories [10, 11].
@@
There are some attempts to unify GUT and family groups into a larger GUT group in 4D and higher-dimensional theories [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. In particular, models based on $SU(19)$ GUT gauge symmetry are discussed in the 4D framework [32, 37] and the six-dimensional (6D) framework by one of the authors [36]. In these models, the three chiral generations of the SM fermions are unified into a fermion multiplet in the second-rank antisymmetric tensor representation ${\bf 171}$ of $SU(19)$. In such models, $SU(19)$ is broken to a maximal regular subgroup $SU(16)\times SU(3)\times U(1)$, and $SU(16)$ is further broken to a maximal special subgroup $SO(10)$. The branching rule of ${\bf 171}$ under $SU(19)\supset SO(10)\times SU(3)\times U(1)$ is given by [39]
@@
(For more information about Lie algebras and their subalgebras, see, e.g., Ref. [39].)
In this paper, we will propose a new model based on $SO(16)$ GUT gauge symmetry that unifies the three chiral generations of SM fermions into a 6D Weyl fermion in a spinor representation of $SO(16)$ without exotic fermions on the 6D spacetime $M^{4}\times D^{2}/\mathbb{Z}_{N}$ $(N\geq 9)$, where $D^{2}$ is a 2-dimensional (2D) disk and there is a fixed point at the origin [42]. The 6D $SO(16)$ gauge anomaly is canceled by the vectorlike nature of the model, and the 4D gauge anomalies are canceled by introducing suitable 4D localized fermions at the fixed point. The gauge coupling constant of $SO(16)$ is asymptotically free.
@@
Let’s consider the 6D spacetime $M^{4}\times D^{2}/\mathbb{Z}_{N}$ with Cartesian coordinates $x^{\mu}$ ($\mu=0,1,2,3$) and circular coordinates $(r,\theta)$ or a complex coordinate $z=re^{i\theta}$, where $D^{2}$ is a 2-dimensional (2D) disk and there is a fixed point at the origin $r=0$ [42]. We take the radius of the disc $D^{2}$ as $R$. A $\mathbb{Z}_{N}$ symmetry on $D^{2}$ is defined as
@@
In the following, we discuss the gauge fields associated with the $SO(16)$ generators $T^{\alpha}$ ($\alpha=1,2,...,120$), which are expressed using the Pauli matrices $\sigma_{a}$ ($a=1,2,3$) and $8\times 8$ symmetric and anti-symmetric matrices $S_{8}$ and $A_{8}$ as [43]:
@@
where $\bar{\omega}:=\omega^{-1}$, $\eta_{\Phi}\subset\mathbb{Z}_{N}$, and $\ell_{\Omega}$ and $m_{\Omega}$ are non-negative integers depending on the representation of $G$. The boundary condition at $r=R$ in $D^{2}/\mathbb{Z}_{N}$ can be chosen as the Dirichlet or Neumann boundary condition. However, if the Dirichlet boundary condition is chosen, there will be no zero mode, so in the ensuing discussion, the Neumann boundary condition will be chosen to allow the existence of zero modes. (For details, see, e.g., Refs. [28, 42].) We will apply this discussion for the $SO(16)$ gauge field $A_{\mu}$ in the 4D spacetime below, noting that different choices of $R_{\Omega}$ will lead to different symmetry-breaking schemes.
@@
This symmetry-breaking pattern is realized by taking the following matrix:
$R_{\Omega}=\mbox{exp}\left[i\frac{2\pi}{N}\sigma_{2}\otimes(-a)I_{8}\right],$ (2.10)
where $a=1,2,\cdots,N-1$ (mod. $N$), and the exponents are proportional to the $U(1)_{a}$ charges. In this case, the gauge field $A_{\mu}$ is transformed under $\mathbb{Z}_{N}$ as
@@
$R_{\Omega}=\mbox{exp}\left[i\frac{2\pi}{N}\sigma_{2}\otimes\left\{\left(-a-3b\right)I_{5}\oplus\left(-a+5b\right)I_{3}\right\}\right],$ (2.17)
where $|a+3b|,|a-5b|\in\mathbb{Z}_{\not=0}$, and the exponents are proportional to the $U(1)_{a}$ and $U(1)_{b}$ charges.
@@
${\bf 120}$ $=({\bf 45,1})(0)\oplus({\bf 1,8})(0)\oplus({\bf 1,1})(0)\oplus({\bf 1,3})(4)\oplus({\bf 1,\overline{3}})(-4)\oplus({\bf 10,3})(-2)\oplus({\bf 10,\overline{3}})(2).$ (2.18)
This symmetry-breaking pattern is realized by choosing $a=-3\ell/4$ and $b=\ell/4$ $(\ell\in\mathbb{Z}_{\not=0})$ in Eq. (2.17) as
$R_{\Omega}=\mbox{exp}\left[i\frac{2\pi}{N}\sigma_{2}\otimes\left\{0\cdot I_{5}\oplus 2\ell\cdot I_{3}\right\}\right],$ (2.19)
where the exponents are proportional to the $U(1)$ charges. This matrix breaks $SO(16)$ to $SO(10)\times SU(3)\times U(1)$. Note that if we choose special values of $N$ and $\ell$, then $SO(16)$ is unbroken for $\ell=0$ (mod. $N/2$), and $SO(16)$ is broken to $SO(10)\times SO(6)\simeq SO(10)\times SU(4)$ for $\ell=N/4$ (mod. $N/2$), where $\ell\in\mathbb{Z}_{\not=0}$. We will consider the $SO(16)$ breaking to $SO(10)\times SU(3)\times U(1)$ in the following sections.
@@ 3 Gauge symmetry and gauge boson
From Table 1, we find that there are the following symmetry breaking patterns, depending on the value of $\ell$:
@@
We focus on the symmetry-breaking pattern $SO(16)\to SO(10)\times SU(3)\times U(1)$ in Eq. (3.5), where the three chiral generations of SM fermions can be unified into one fermion multiplet. In this case, we can take $\ell\not=0$ (mod. $N/4$). Here, we further impose the condition that no extra zero mode appears from the extra-dimensional components of the gauge field. From Table 1, we find the following constraints for $\ell$ and $N$ to avoid the unwanted zero modes:
@@
This leads to the requirement that $N\geq 6$. For $\ell=2$, from Eq. (3.6), we find
@@
This leads to $N=6$ or $N\geq 10$. We can also consider $\ell\geq 3$ and negative integers in the same manner.
@@ 4 Unification of the three chiral generations of SM fermions
$\Omega\Psi_{\bf 128}^{\alpha}(x^{\mu},z,\bar{z})=\Psi_{\bf 128}^{\alpha}(x^{\mu},\omega z,\bar{\omega}\bar{z})=\eta_{\Psi^{\alpha}}R_{\Omega}^{\rm sp}\Psi_{\bf 128}^{\alpha}(x^{\mu},z,\bar{z}),$ (4.4)
@@
We need to find a set of $\ell$, $k$ and $N$ that simultaneously satisfy Eqs. (4.5) and (4.6). There turns out to be an infinite number of allowed sets of $\ell$, $k$ and $N$. Hence, we will only consider a few simple examples. First, we set $\ell=1$. In this case, Eq. (4.5) gives
@@
This leads to the requirement $N\geq 9$. In other words, we have now found a possible set of parameters, $(\ell,k,N)=(1,-4,9)$.
@@
This then leads to $N=4$ or $N\geq 8$. Combining the two conditions, we need to take $N\geq 10$. Thus, we have found another possible set of parameters: $(\ell,k,N)=(2,0,10)$.
In summary, to unify the three chiral generations of SM fermions without the occurrence of exotic fermions, one should take $(N,\eta_{\Psi^{1}},\eta_{\Psi^{2}})=(9,\omega^{-1},\omega^{-4})$, $(10,\omega^{-2},1)$, $\dots$ In such cases, only the fermions in the $({\bf 16,3})(1)$ representation of $SO(10)\times SU(3)\times U(1)$ have zero modes.
@@
From the above discussions, it is clear that one choice of having the three chiral generations of SM fermions is to take $(N,\eta_{\Psi^{1}},\eta_{\Psi^{2}})=(9,\omega^{-1},\omega^{-4})$. In this case, we obtain the 4D left-handed fermions in $({\bf 16,3})(1)$ of $SO(10)\times SU(3)\times U(1)$ from a single 6D fermion $\Psi_{\bf 128}^{\alpha=1}$ without having exotic fermions. In the following, we discuss 6D and 4D gauge anomaly cancellation in this scenario.
@@
$A_{\rm bulk}^{U(1)\mbox{-}U(1)\mbox{-}U(1)}=16\times 3\times(+1)^{3}=+48,$
@@
$A_{\rm bulk}^{\mbox{grav}\mbox{-}\mbox{grav}\mbox{-}U(1)}=16\times 3\times T_{SO(4)}{(\bf 2,1)}\times(+1)=+24,$ (5.1)
@@
We now examine what kind of 4D fields should be introduced to cancel out the gauge anomaly generated by the 6D fields given in Eq. (5.1). It turns out to be impossible to uniquely determine the required set of 4D fields solely by the requirement that the 4D gauge anomalies arising from the 6D fields cancel out. In the following, we provide an explicit example. One way is to introduce 4D localized left-handed fermions in $({\bf 10,\overline{3}})(-1)$ and $({\bf 1,3})(+1)$ of $SO(10)\times SU(3)\times U(1)$, denoted by $\psi_{L({\bf 10,\overline{3}})(-1)}^{(\beta)}$ $(\beta=1,2)$ and $\psi_{L({\bf 1,{3}})(+1)}^{(\gamma)}$ $(\gamma=1,2,3,4)$, respectively. In the case, the $SO(10)\mbox{-}SO(10)\mbox{-}U(1)$ mixed gauge anomaly is canceled by $\psi_{L({\bf 10,\overline{3}})(-1)}^{(\beta)}$; the remaining anomalies are canceled by the combination of $\psi_{L({\bf 10,\overline{3}})(-1)}^{(\beta)}$ and $\psi_{L({\bf 1,{3}})(+1)}^{(\gamma)}$. When we take into account the vacuum expectation values (VEVs) of the $SU(3)$- and $U(1)$-breaking scalar fields, these 4D fields obtain masses that do not appear at low energies and will not be discussed further in this paper.
@@ 6 Symmetry breaking via the VEVs of scalars
Next, we consider the scalar content to realize the SM gauge symmetry at low energies. The set of scalar fields, which properly break the $SO(16)$ symmetry into $G_{\rm SM}$, is also not uniquely determined. Here we provide a working example by introducing localized scalars $\phi_{{\bf(10,1)}(0)}$, $\phi_{{\bf(45,1})(0)}$, $\phi_{{\bf(16,1)}(0)}$, $\phi_{{\bf(1,3)}(0)}$, $\phi_{{\bf(1,6)}(+2)}$, and $\phi_{{\bf(1,1)}(+1)}$ to realize $G_{\rm SM}$ at low energies. In this scenario, the symmetry-breaking pattern is given as follows:
$\begin{array}[]{rll}SO(16)\ \to&SO(10)\times SU(3)\times U(1)&\mbox{by the non-trivial space structure}\\ \to&SO(10)&\mbox{by the VEVs}\ \langle\phi_{{\bf(1,3)}(0)}\rangle,\langle\phi_{{\bf(1,6)}(+2)}\rangle,\langle\phi_{{\bf(1,1)}(+1)}\rangle\not=0\\ \to&G_{\rm SM}&\mbox{by the VEVs}\ \langle\phi_{{\bf(45,1)}(0)}\rangle,\langle\phi_{{\bf(16,1)}(0)}\rangle\not=0\\ \to&SU(3)_{C}\times U(1)_{\rm EM}&\mbox{by the VEV}\ \langle\phi_{{\bf(10,1)}(0)}\rangle\not=0,\\ \end{array}$ (6.5)
where we assume that the Kaluza-Klein (KK) mass scales $1/R$ and the non-vanishing VEVs are at the order of $M_{\rm GUT}\simeq O(10^{16}-10^{17})$ GeV except the VEV of the SM Higgs $\langle\phi_{{\bf(10,1)}(0)}\rangle$. By assuming symmetry breaking with the scalar field VEVs given above, all the introduced 4D fermions will attain large mass and will not appear at low energies. The $SO(10)$ GUT symmetry breaking to the SM gauge symmetry has been thoroughly examined in, for example, Refs. [47, 48, 49, 50, 39] and will not be discussed further here.